# Terminal 3-folds that are not Cohen-Macaulay

### v1

@@
An important ingredient of the minimal model program is that Kawamata log terminal singularities in characteristic zero are rational, and in particular Cohen-Macaulay. In the special case of cone singularities, this fact is related to the Kodaira vanishing theorem restricted to Fano varieties. It turns out that Kodaira vanishing fails even for Fano varieties, in every characteristic $p>0$ [26]. This led to examples of klt, and even terminal, singularities that are not Cohen-Macaulay [26, 28]. (Terminal singularities are the smallest class of singularities that can be allowed on minimal models.)
@@
Then $X$ is terminal but not Cohen-Macaulay [26, Theorem 5.1]. This is the lowest possible dimension, because every terminal (or just normal) surface is Cohen-Macaulay. Cohen-Macaulayness and stronger properties such as $F$-regularity help to construct contractions of varieties. Partly for this reason, the MMP for 3-folds is known only in characteristics at least 5 [13]. By Arvidsson–Bernasconi–Lacini, klt singularities in characteristic greater than 5 are Cohen-Macaulay, whereas there are klt singularities that are not Cohen-Macaulay in characteristics 2, 3, and 5 [4, 6, 9].
In this paper, we construct terminal 3-fold singularities that are not Cohen-Macaulay in five new cases: mixed characteristic $(0,2)$, characteristic 3, mixed characteristic $(0,3)$, characteristic 5, and mixed characteristic $(0,5)$ (Theorems 0.1, 6.1, 7.1, 8.1, and 9.1. This is optimal, in view of the result of Arvidsson–Bernasconi–Lacini. Indeed, the MMP for schemes of dimension 3 was developed in mixed characteristic when the residue characteristic is greater than 5 [7, 25]. This raised the question of whether vanishing theorems for 3-folds hold in mixed characteristic. Given our counterexample over $\text{\bf F}_{2}$, one might expect an example of dimension 4, flat over the 2-adic integers $\text{\bf Z}_{2}$, with fiber over $\text{\bf F}_{2}$ being the 3-fold singularity above. In fact, each of our examples has dimension 3 as a scheme. For example, over $\text{\bf Z}_{2}$ we have:
@@
Note that an action of a $p$-group on a smooth variety in characteristic $p$ with an isolated fixed point is never formally isomorphic to a linear action, since a nonzero representation $V$ of a $p$-group $G$ in characteristic $p$ has nonzero $G$-fixed subspace. In fact, there are continuous families of non-equivalent actions on smooth varieties in characteristic $p$, and likewise on regular schemes of mixed characteristic. For an action of $G=\text{\bf Z}/p$ on a regular scheme $W$ of dimension 3 with an isolated fixed point in characteristic $p$ (as in Theorem 0.1), it is common for $W/G$ not to be Cohen-Macaulay, essentially because the cohomology of $G$ contributes to the local cohomology of $W/G$. The difficulty is to construct an example with $W/G$ terminal. For a more complicated action of $G$, the quotient scheme would usually not be terminal or even log canonical. To find the examples in this paper, the idea was to look for the simplest possible actions of $G=\text{\bf Z}/p$ on a regular 3-dimensional scheme with an isolated fixed point of residue characteristic $p$.
@@
These examples should lead to other failures of vanishing theorems. In particular, by Baudin, Bernasconi, and Kawakami, these examples imply that Frobenius-stable Grauert-Riemenschneider vanishing fails in characteristic 2, 3, and 5 [5, Remark 1.2].
@@ 1 Notation
We write $R\{x_{1},\ldots,x_{n}\}$ to mean the free module over a ring $R$ with basis elements $x_{1},\ldots,x_{n}$.
@@
See section 5 for the definition of terminal singularities on general schemes, following [18, section 2.1].
For a positive integer $r$, let $\mu_{r}$ be the group scheme (over any base scheme) of $r$th roots of unity. The Reid-Tai criterion is the following description of which cyclic quotient singularities are canonical or terminal [23, Theorem 4.11]. This is often stated over a field, but it works even in mixed characteristic for $\mu_{r}$-quotient singularities. The point is that Kato’s theory of log regular schemes provides a mixed-characteristic analog of toric singularities, which includes the case of $\mu_{r}$-quotient singularities [15]. For such schemes, resolutions of singularities and the canonical divisor can be described in purely combinatorial terms. (The proof of Theorem 5.1 explains how to interpret the canonical divisor and terminal singularities in mixed characteristic.) In the following criterion, for $a$ an integer and $b$ a positive integer, consider $a\bmod b$ as an integer in the set $\{0,\ldots,b-1\}$.
@@
Under the assumptions of Theorem 1.1, let $t_{1},\ldots,t_{n}$ be coordinates for $X$ at $P$ that are $\mu_{r}$-eigenfunctions with weights $b_{1},\ldots,b_{n}$. For $i=1,\ldots,n$, let $D_{i}$ be the irreducible divisor on $X/\mu_{r}$ that is the image of $\{t_{i}=0\}$ in $X$. Let $c_{1},\ldots,c_{n}$ be real numbers. Then the pair $(X,\sum c_{j}D_{j})$ is terminal near the image of $P$ if and only if $c_{j}<1$ for each $j$, $c_{j}+c_{k}<1$ for each $j\neq k$, and
@@ 2 Recognizing $\mu_{p}$-quotient singularities
Király and Lütkebohmert analyzed when a quotient scheme by $\text{\bf Z}/p$ is regular, the hard case being where the residue characteristic is $p$ [16]. More generally, we now give a sufficient condition for a quotient by $\text{\bf Z}/p$ to be a toric singularity, or in particular when it is isomorphic to a $\mu_{p}$-quotient singularity. (Actions of $\mu_{p}$ are far simpler than actions of $\text{\bf Z}/p$: they are linearizable near a fixed point, like finite group actions in characteristic zero, because $\mu_{p}$ is linearly reductive. Also, $\mu_{p}$-quotient singularities are always Cohen-Macaulay.)
@@
Here is Király and Lütkebohmert’s main result [16, Theorem 2]. (See Theorem 2.6 for a more detailed statement.)
@@
One might ask whether the assumption that $p\in e^{p-1}\operatorname{\mathfrak{m}}$ in Theorem 2.2 can be omitted. Fortunately, this assumption is automatic in characteristic $p$, and it will be easy to check in our mixed characteristic examples.
@@
The fixed point scheme $U^{G}$ is defined as the closed subscheme of $U$ cut out by the ideal generated by $I(O(U))$. The following formulas hold for any action of $G$ on a commutative ring [16, Remark 3]:
@@
We are given coordinates $x_{1},\ldots,x_{n}$ for $U$ near $P$. Under our assumptions ($U$ regular and $k_{U/G}=k_{U}$), the ideal generated by $I(O_{U,P})$ in $O_{U,P}$ is generated by $I(x_{1}),\ldots,I(x_{n})$ [16, Proposition 6]. That is, after shrinking $U$ around $P$ if necessary, the fixed point scheme $U_{G}$ is the closed subscheme defined by $I(x_{1}),\ldots,I(x_{n})$.
@@
Since the ring $O_{U,P}$ is regular, it is a domain, and so $I(y)/e$ is well-defined for each element $y$ in $\operatorname{\mathfrak{m}}$. Since $G$ fixes the point $P$, $G$ maps $\operatorname{\mathfrak{m}}$ into itself. Since $I(\operatorname{\mathfrak{m}})\subset e\operatorname{\mathfrak{m}}$, Lemma 2.3 gives that $I(\operatorname{\mathfrak{m}}^{2})\subset e\operatorname{\mathfrak{m}}^{2}$. Since $I$ is additive, it follows that $\varphi$ is a well-defined additive function from $\operatorname{\mathfrak{m}}/\operatorname{\mathfrak{m}}^{2}$ to itself. By Lemma 2.3, $\varphi$ is linear over the ring of invariants $(O_{U,P})^{G}$. Since that ring has residue field $k_{U/G}=k_{U}$, $\varphi$ is $k_{U}$-linear. ∎
@@
The statement is clear if $\zeta=1$. So assume that $\zeta\neq 1$. Since $O(U)$ is a domain and $(1-\zeta)(1+\zeta+\cdots+\zeta^{p-1})=1-\zeta^{p}=0$, we must have $01+\zeta+\cdots+\zeta^{p-1}=$ 0. That is, we have a homomorphism from the ring of integers $\text{\bf Z}[\mu_{p}]$ in $\text{\bf Q}(\mu_{p})$ to $O(U)$, taking a primitive $p$th root of unity to $\zeta$. In $\text{\bf Z}[\mu_{p}]$, $p$ is $(1-\zeta)^{p-1}$ times a unit [19, section IV.1], and so the same is true in $O(U)$. We are given that $p$ is in $e^{p-1}\operatorname{\mathfrak{m}}$. Since $U$ is regular, $O_{U,P}$ is a unique factorization domain, and so $1-\zeta$ must be a multiple of $e$; write $1-\zeta=ea$ for some $a\in O_{U,P}$. If $a$ is a unit, then $p$ would be $e^{p-1}$ times a unit, a contradiction. So, as an element of $O_{U,P}$, $1-\zeta$ is in $e\operatorname{\mathfrak{m}}$. After shrinking $U$ if necessary, this gives the same conclusion in $O(U)$. ∎
@@
We need to show that $f$ is not a $p$th power in $O(U)$. Suppose it is, say $f=u^{p}$ for a regular function $u$ on $U$. Since $f$ is a unit, so is $u$. Then $1/g^{p}=\sigma(f)/f=(\sigma(u)/u)^{p}$, and so $\zeta/g=\sigma(u)/u$ for some $p$th root of unity $\zeta$ in $O(U)$. Here $\zeta=1+(\zeta-1)=1+e\operatorname{\mathfrak{m}}$, by Lemma 2.5, and $1/g=1-e(1+\operatorname{\mathfrak{m}})$, so $\sigma(u)/u=1-e(1+\operatorname{\mathfrak{m}})$. Here $\sigma(u)/u=1+I(u)/u$, so $I(u)=ue(1+\operatorname{\mathfrak{m}})=e(\text{unit})$. This contradicts our assumption that $I(O(U))$ is contained in $e\operatorname{\mathfrak{m}}$. So in fact $f$ is not a $p$th power.
From there, we can show that the scheme $W$ is integral (after shrinking $U$ near $P$, if necessary). Namely, since $f$ is not a $p$th power in $O(U)$, $f$ is also not a $p$th power in the function field $k(U)$, and so $k(U)[f^{1/p}]$ is a degree-$p$ field extension of $k(U)$. Write $\alpha$ for the $\mu_{p}$-torsor $W\to U$. Since $W=\{w^{p}=f\}$, there is a nonempty open subset $V\subset U$ with $\alpha^{-1}(V)$ integral. Since $W\to U$ is finite and flat, $W$ is Cohen-Macaulay and equidimensional. By equidimensionality, every irreducible component of $W$ must dominate $U$. Since $\alpha^{-1}(V)$ is irreducible, it follows that $W$ is irreducible. Since $\alpha^{-1}(V)$ is reduced, $W$ is reduced in codimension 0; since $W$ is Cohen-Macaulay, it follows that $W$ is reduced [24, Tag 031R]. Since $W$ is reduced and irreducible, it is integral.
@@
Statement (1) is due to Király and Lütkebohmert for $B$ normal, but their proof works without change for $B$ a domain [16, Theorem 2 and Proposition 5]. They also prove statement (2). They prove statement (3) only when $B$ is regular. We now extend the proof of (3) for $B$ only a domain.
@@
Let us write out the action of $G$ on $W$. The maximal ideal of $P$ in $W$ is generated by $w_{0},x_{1},\ldots,x_{n}$. We have $I(w_{0})=I(w)=w(\frac{1}{g}-1)=(1+w_{0})(\frac{1}{g}-1)$. Since $1/g=1+eu$ for some unit $u$ on $U$, we have $I(w_{0})=eu(1+w_{0})=e(\text{unit})$. We also have $I(x_{i})$ in the ideal $(e)$ for $i=1,\ldots,n$; so the fixed point scheme $W^{G}$ is defined by the single equation $e=0$ in $W$. As a result, even though $W$ is typically not normal, Theorem 2.6 gives that the morphism $W\to W/G=Q$ is finite and faithfully flat of degree $p$. It follows that $Q$ is noetherian [24, Tag 033E]. (Beware that for a general noetherian scheme $X$ with an action of a finite group $G$, $X/G$ need not be noetherian, and the morphism $X\to X/G$ need not be finite [22, Proposition 0.10]. These properties do hold if $X$ is of finite type over a noetherian ring $A$ and $G$ acts $A$-linearly [11, Theorem and Corollary 4].)
The action of $\mu_{p}$ on the affine scheme $Q$ gives a grading of $O(Q)$ by $\text{\bf Z}/p$. For each $j\in\text{\bf Z}/p$, since $O(Q)$ is noetherian, the ideal in $O(Q)$ generated by the $j$th graded piece $O(Q)_{j}$ is finitely generated, and so $O(Q)_{j}$ is a finitely generated module over $O(Q)_{0}=O(Q/\mu_{p})$. So the whole ring $O(Q)$ is finite over $O(Q/\mu_{p})$; that is, $Q\to Q/\mu_{p}$ is finite. Also, $O(Q/\mu_{p})$ is a pure subring (because it is a summand) of the noetherian ring $O(Q)$, so it is noetherian [14, Proposition 6.15]; that is, $Q/\mu_{p}=U/G$ is noetherian. Finally, the composition $W\to Q\to Q/\mu_{p}=U/G$ is finite, and $O(U)$ is a sub-$O(U/G)$-module of $O(W)$, so $O(U)$ is a finitely generated $O(U/G)$-module; that is, $U\to U/G$ is finite.
@@
The number $r$ of generators $h_{1},\ldots,h_{r}$ for the maximal ideal $\operatorname{\mathfrak{m}}_{Q}$ in the local ring $O_{Q,P}$ is at least $n=\dim(Q)$. On the other hand, Lemma 2.7 implies that the extended ideal $(h_{1},\ldots,h_{r})$ in $O_{W,P}$ can be generated by only $n$ elements, so the vector space $(h_{1},\ldots,h_{r})\otimes_{O_{W,P}}k_{W}$ has dimension at most $n$. So this vector space is spanned by $n$ of the $h_{i}$’s, which we can assume are $h_{1},\ldots,h_{n}$. By Nakayama’s lemma, it follows that the extended ideal $(h_{1},\ldots,h_{r})$ is equal to the extended ideal $(h_{1},\ldots,h_{n})$ in $O_{W,P}$. Since $W\to Q$ is faithfully flat, extending and contracting an ideal in $O_{Q,P}$ gives the same ideal [20, Theorem 7.5]. As a result, we have $(h_{1},\ldots,h_{n})=(h_{1},\ldots,h_{r})$ in $O_{Q,P}$. That is, the maximal ideal $\operatorname{\mathfrak{m}}_{Q}$ can be generated by $n=\dim(Q)$ elements, which means that $Q$ is regular. (This is somewhat surprising, since $W$ is typically not regular or even normal.)
@@
using that $e$ is in $\operatorname{\mathfrak{m}}_{U}$. Since the function $g_{i}w^{b_{i}}$ is $G$-invariant on $W$, we have $0=I(g_{i}w^{b_{i}})=\sigma(g_{i})I(w^{b_{i}})+I(g_{i})w^{b_{i}}=(g_{i}+I(g_{i}))I(w^{b_{i}})+I(g_{i})w^{b_{i}}$. We want to consider this equality modulo $e\operatorname{\mathfrak{m}}_{U}^{2}O(W)$, and so the term $I(g_{i})I(w^{b_{i}})$ can be omitted. Namely, we have

### v2

@@
An important ingredient of the minimal model program is that Kawamata log terminal singularities in characteristic zero are rational, and in particular Cohen-Macaulay. In the special case of cone singularities, this fact is related to the Kodaira vanishing theorem restricted to Fano varieties. It turns out that Kodaira vanishing fails even for Fano varieties, in every characteristic $p>0$ [29]. This led to examples of klt, and even terminal, singularities that are not Cohen-Macaulay [21, 29, 31]. (Terminal singularities are the smallest class of singularities that can be allowed on minimal models.)
@@
Then $X$ is terminal but not Cohen-Macaulay [29, Theorem 5.1]. This is the lowest possible dimension, because every terminal (or just normal) surface is Cohen-Macaulay. Cohen-Macaulayness and stronger properties such as $F$-regularity help to construct contractions of varieties. Partly for this reason, the MMP for 3-folds is known only in characteristics at least 5 [15, 8, 14]. By Arvidsson–Bernasconi–Lacini, klt singularities in characteristic greater than 5 are Cohen-Macaulay, whereas there are klt singularities that are not Cohen-Macaulay in characteristics 2, 3, and 5 [4, 6, 10].
In this paper, we construct terminal 3-fold singularities that are not Cohen-Macaulay in five new cases: mixed characteristic $(0,2)$, characteristic 3, mixed characteristic $(0,3)$, characteristic 5, and mixed characteristic $(0,5)$ (Theorems 0.1, 6.1, 7.1, 8.1, and 9.1). This is optimal, in view of the result of Arvidsson–Bernasconi–Lacini. Indeed, the MMP for schemes of dimension 3 was developed in mixed characteristic when the residue characteristic is greater than 5 [7, 28]. This raised the question of whether vanishing theorems for 3-folds hold in mixed characteristic. Given our counterexample over $\text{\bf F}_{2}$, one might expect an example of dimension 4, flat over the 2-adic integers $\text{\bf Z}_{2}$, with fiber over $\text{\bf F}_{2}$ being the 3-fold singularity above. In fact, each of our examples has dimension 3 as a scheme. For example, over $\text{\bf Z}_{2}$ we have:
@@
Note that an action of a $p$-group with an isolated fixed point on a positive-dimensional smooth variety in characteristic $p$ is never formally isomorphic to a linear action, because a nonzero representation $V$ of a $p$-group $G$ in characteristic $p$ has nonzero $G$-fixed subspace. In fact, there are continuous families of non-equivalent actions on smooth varieties in characteristic $p$, and likewise on regular schemes of mixed characteristic. For an action of $G=\text{\bf Z}/p$ on a regular scheme $W$ of dimension 3 with an isolated fixed point in characteristic $p$ (as in Theorem 0.1), it is common for $W/G$ not to be Cohen-Macaulay, essentially because the cohomology of $G$ contributes to the local cohomology of $W/G$. The difficulty is to construct an example with $W/G$ terminal. For a more complicated action of $G$, the quotient scheme would usually not be terminal or even log canonical. To find the examples in this paper, the idea was to look for the simplest possible actions of $G=\text{\bf Z}/p$ on a regular 3-dimensional scheme with an isolated fixed point of residue characteristic $p$.
@@
These examples should lead to other failures of vanishing theorems. In particular, by Baudin, Bernasconi, and Kawakami, these examples imply that Frobenius-stable Grauert-Riemenschneider vanishing fails for terminal 3-folds in characteristic 2, 3, or 5 [5, Remark 1.2].
@@ 1 Notation
We write $R\{x_{1},\ldots,x_{n}\}$ for the free module over a ring $R$ with basis elements $x_{1},\ldots,x_{n}$.
@@
See section 5 for the definition of the canonical divisor and terminal singularities on general schemes, following [20, section 2.1].
For a positive integer $r$, let $\mu_{r}$ be the group scheme (over any base scheme) of $r$th roots of unity. The Reid-Tai criterion is the following description of which cyclic quotient singularities are canonical or terminal [25, Theorem 4.11]. This is often stated over a field, but it works even in mixed characteristic for $\mu_{r}$-quotient singularities. The point is that Kato’s theory of log regular schemes provides a mixed-characteristic analog of toric singularities, which includes the case of $\mu_{r}$-quotient singularities [17]. For such schemes, resolutions of singularities and the canonical divisor can be described in purely combinatorial terms. In the following criterion, for $a$ an integer and $b$ a positive integer, consider $a\bmod b$ as an integer in the set $\{0,\ldots,b-1\}$.
@@
Under the assumptions of Theorem 1.1, let $t_{1},\ldots,t_{n}$ be coordinates for $X$ at $P$ that are $\mu_{r}$-eigenfunctions with weights $b_{1},\ldots,b_{n}$. For $i=1,\ldots,n$, let $D_{i}$ be the irreducible divisor on $X/\mu_{r}$ that is the image of $\{t_{i}=0\}$ in $X$. Let $c_{1},\ldots,c_{n}$ be real numbers. Then the pair $(X/\mu_{r},\sum c_{j}D_{j})$ is terminal near the image of $P$ if and only if $c_{j}<1$ for each $j$, $c_{j}+c_{k}<1$ for each $j\neq k$, and
@@ 2 Recognizing $\mu_{p}$-quotient singularities
Király and Lütkebohmert analyzed when a quotient scheme by $\text{\bf Z}/p$ is regular, the hard case being where the residue characteristic is $p$ [18]. More generally, we now give a sufficient condition for a quotient by $\text{\bf Z}/p$ to be a toric singularity, or in particular to be a $\mu_{p}$-quotient singularity. (Actions of $\mu_{p}$ are far simpler than actions of $\text{\bf Z}/p$: they are linearizable near a fixed point, like finite group actions in characteristic zero, because $\mu_{p}$ is linearly reductive. Also, $\mu_{p}$-quotient singularities are always Cohen-Macaulay.)
@@
Here is Király and Lütkebohmert’s main result [18, Theorem 2]. (See Theorem 2.6 for a more detailed statement.)
@@
One might ask whether the assumption that $p\in e^{p-1}\operatorname{\mathfrak{m}}$ in Theorem 2.2 can be omitted. Fortunately, this assumption is automatic in characteristic $p$, and it will be easy to check in our mixed-characteristic examples.
@@
The fixed point scheme $U^{G}$ is defined as the closed subscheme of $U$ cut out by the ideal generated by $I(O(U))$. The following formulas hold for any action of $G$ on a commutative ring [18, Remark 3]:
@@
We are given coordinates $x_{1},\ldots,x_{n}$ for $U$ near $P$. Under our assumptions ($U$ regular and $k_{U/G}=k_{U}$), the ideal generated by $I(O_{U,P})$ in $O_{U,P}$ is generated by $I(x_{1}),\ldots,I(x_{n})$ [18, Proposition 6]. That is, after shrinking $U$ around $P$ if necessary, the fixed point scheme $U^{G}$ is the closed subscheme defined by $I(x_{1}),\ldots,I(x_{n})$.
@@
Since the local ring $O_{U,P}$ is regular, it is a domain, and so $I(y)/e$ is well-defined for each element $y$ in $\operatorname{\mathfrak{m}}$. Since $G$ fixes the point $P$, $G$ maps $\operatorname{\mathfrak{m}}$ into itself. Since $I(\operatorname{\mathfrak{m}})\subset e\operatorname{\mathfrak{m}}$, Lemma 2.3 gives that $I(\operatorname{\mathfrak{m}}^{2})\subset e\operatorname{\mathfrak{m}}^{2}$. Since $I$ is additive, it follows that $\varphi$ is a well-defined additive function from $\operatorname{\mathfrak{m}}/\operatorname{\mathfrak{m}}^{2}$ to itself. By Lemma 2.3, $\varphi$ is linear over the ring of invariants $(O_{U,P})^{G}$. Since that ring has residue field $k_{U/G}=k_{U}$, $\varphi$ is $k_{U}$-linear. ∎
@@
The statement is clear if $\zeta=1$. So assume that $\zeta\neq 1$. Since $O(U)$ is a domain and $(1-\zeta)(1+\zeta+\cdots+\zeta^{p-1})=1-\zeta^{p}=0$, we must have $01+\zeta+\cdots+\zeta^{p-1}=$ 0. That is, we have a homomorphism from the ring of integers $\text{\bf Z}[\mu_{p}]$ in $\text{\bf Q}(\mu_{p})$ to $O(U)$, taking a primitive $p$th root of unity to $\zeta$. In $\text{\bf Z}[\mu_{p}]$, $p$ is $(1-\zeta)^{p-1}$ times a unit [22, section IV.1], and so the same is true in $O(U)$. We are given that $p$ is in $e^{p-1}\operatorname{\mathfrak{m}}$. Since $U$ is regular, $O_{U,P}$ is a unique factorization domain, and so $1-\zeta$ must be a multiple of $e$; write $1-\zeta=ea$ for some $a\in O_{U,P}$. If $a$ is a unit, then $p$ would be $e^{p-1}$ times a unit, a contradiction. So, as an element of $O_{U,P}$, $1-\zeta$ is in $e\operatorname{\mathfrak{m}}$. After shrinking $U$ around $P$ if necessary, this gives the same conclusion in $O(U)$. ∎
@@
We need to show that $f$ is not a $p$th power in $O(U)$. Suppose it is, say $f=u^{p}$ for a regular function $u$ on $U$. Since $f$ is a unit, so is $u$. Then $1/g^{p}=\sigma(f)/f=(\sigma(u)/u)^{p}$, and so $\zeta/g=\sigma(u)/u$ for some $p$th root of unity $\zeta$ in $O(U)$. Here $\zeta=1+(\zeta-1)=1+e\operatorname{\mathfrak{m}}$ by Lemma 2.5, and $1/g=1-e(1+\operatorname{\mathfrak{m}})$, so $\sigma(u)/u=1-e(1+\operatorname{\mathfrak{m}})$. Here $\sigma(u)/u=1+I(u)/u$, so $I(u)=-ue(1+\operatorname{\mathfrak{m}})=e(\text{unit})$. This contradicts our assumption that $I(O(U))$ is contained in $e\operatorname{\mathfrak{m}}$. So in fact $f$ is not a $p$th power.
From there, we can show that the scheme $W$ is integral (after shrinking $U$ around $P$, if necessary). Namely, since $f$ is not a $p$th power in $O(U)$, $f$ is also not a $p$th power in the function field $k(U)$, and so $k(U)[f^{1/p}]$ is a degree-$p$ field extension of $k(U)$. Write $\alpha$ for the $\mu_{p}$-torsor $W\to U$. Since $W=\{w^{p}=f\}$, there is a nonempty open subset $V\subset U$ with $\alpha^{-1}(V)$ integral. Since $W\to U$ is finite and flat, $W$ is Cohen-Macaulay and equidimensional. By equidimensionality, every irreducible component of $W$ must dominate $U$. Since $\alpha^{-1}(V)$ is irreducible, it follows that $W$ is irreducible. Since $\alpha^{-1}(V)$ is reduced, $W$ is reduced in codimension 0; since $W$ is Cohen-Macaulay, it follows that $W$ is reduced [27, Tag 031R]. Since $W$ is reduced and irreducible, it is integral.
@@
Statement (1) is due to Király and Lütkebohmert for $B$ normal, but their proof works without change for $B$ a domain [18, Theorem 2 and Proposition 5]. They also prove statement (2). They prove statement (3) when $B$ is regular. We now extend the proof of (3) for $B$ only a domain.
@@
Let us write out the action of $G$ on $W$. The maximal ideal of $P$ in $W$ is generated by $w_{0},x_{1},\ldots,x_{n}$. We have $I(w_{0})=I(w)=w(\frac{1}{g}-1)=(1+w_{0})(\frac{1}{g}-1)$. Since $1/g=1+eu$ for some unit $u$ on $U$, we have $I(w_{0})=eu(1+w_{0})=e(\text{unit})$. We also have $I(x_{i})$ in the ideal $(e)$ for $i=1,\ldots,n$; so the fixed point scheme $W^{G}$ is defined by the single equation $e=0$ in $W$. As a result, even though $W$ is typically not normal, Theorem 2.6 gives that the morphism $W\to W/G=Q$ is finite and faithfully flat of degree $p$. It follows that $Q$ is noetherian [27, Tag 033E]. (Beware that for a general noetherian scheme $X$ with an action of a finite group $G$, $X/G$ need not be noetherian, and the morphism $X\to X/G$ need not be finite [24, Proposition 0.10]. These properties do hold if $X$ is of finite type over a noetherian ring $A$ and $G$ acts $A$-linearly [12, Theorem and Corollary 4].)
The action of $\mu_{p}$ on the affine scheme $Q$ gives a grading of $O(Q)$ by $\text{\bf Z}/p$. For each $j\in\text{\bf Z}/p$, since $O(Q)$ is noetherian, the ideal in $O(Q)$ generated by the $j$th graded piece $O(Q)_{j}$ is finitely generated, and so $O(Q)_{j}$ is a finitely generated module over $O(Q)_{0}=O(Q/\mu_{p})$. So the whole ring $O(Q)$ is finite over $O(Q/\mu_{p})$; that is, $Q\to Q/\mu_{p}$ is finite. Also, $O(Q/\mu_{p})$ is a pure subring (because it is a summand) of the noetherian ring $O(Q)$, so it is noetherian [16, Proposition 6.15]; that is, $Q/\mu_{p}=U/G$ is noetherian. Finally, the composition $W\to Q\to Q/\mu_{p}=U/G$ is finite, and $O(U)$ is a sub-$O(U/G)$-module of $O(W)$, so $O(U)$ is a finitely generated $O(U/G)$-module; that is, $U\to U/G$ is finite.
@@
The number $r$ of generators $h_{1},\ldots,h_{r}$ for the maximal ideal $\operatorname{\mathfrak{m}}_{Q}$ in the local ring $O_{Q,P}$ is at least $n=\dim(Q)$. On the other hand, Lemma 2.7 implies that the extended ideal $(h_{1},\ldots,h_{r})$ in $O_{W,P}$ can be generated by only $n$ elements, so the vector space $(h_{1},\ldots,h_{r})\otimes_{O_{W,P}}k_{W}$ has dimension at most $n$. So this vector space is spanned by $n$ of the $h_{i}$’s, which we can assume are $h_{1},\ldots,h_{n}$. By Nakayama’s lemma, it follows that the extended ideal $(h_{1},\ldots,h_{r})$ is equal to the extended ideal $(h_{1},\ldots,h_{n})$ in $O_{W,P}$. Since $W\to Q$ is faithfully flat, extending and contracting an ideal in $O_{Q,P}$ gives the same ideal [27, Tag 05CK]. As a result, we have $(h_{1},\ldots,h_{n})=(h_{1},\ldots,h_{r})$ in $O_{Q,P}$. That is, the maximal ideal $\operatorname{\mathfrak{m}}_{Q}$ can be generated by $n=\dim(Q)$ elements, which means that $Q$ is regular. (This is somewhat surprising, since $W$ is typically not regular or even normal.)
@@
using that $e$ is in $\operatorname{\mathfrak{m}}_{U}$. Since the function $g_{i}w^{b_{i}}$ is $G$-invariant on $W$, we have $0=I(g_{i}w^{b_{i}})=\sigma(g_{i})I(w^{b_{i}})+I(g_{i})w^{b_{i}}=(g_{i}+I(g_{i}))I(w^{b_{i}})+I(g_{i})w^{b_{i}}$. When we consider this equality modulo $e\operatorname{\mathfrak{m}}_{U}^{2}O(W)$, the term $I(g_{i})I(w^{b_{i}})$ can be omitted. Namely, we have